The sum of two numbers is 30. The first number is twice the second number. What are the numbers?

Answer by umail08(87)

x = second number
2x = first number
2x + x = 30
3x = 30
3x/3 = 30/3
x = 10
Therefore 2x = 2(10) = 20
So the numbers are 10 and 20